Consider two different implementations of the same instruction set architecture. There are 5 classes of instructions, A, B, C, D, and E. The clock rate and CPI of each class is given in the following table.

Machine

Clock Rate

CPI class A

CPI class B

CPI class C

CPI class D

CPI class E

P1

1.0 GHz

1

1

2

3

2

P2

1.5 GHz

1

2

3

4

3

  1. What are the performances of P1 and P2 expressed in instructions per second?

  2. The number of instructions for classes A, B, C, D is divided equally except for Class E which occurs three times as often as each of the others. Find the execution time on machines P1, P2. Find the performance of the machines.
Answer #1

Performance = clock rate / CPI

P1 ( using CPI class A) = 1.0 * 109 / 1 = 1*109 inst/s

P2 ( using CPI class A) = 1.5 * 109 / 1 = 1.5*109 inst/s

Execution time = cycles / Clock rate

P1 cycles = 1 + 1 + 2 + 3 + ( 2*3) = 13

Execution time on P1 = 13 / 1 = 13 ns

P2 cycles = 1 + 2 + 3 + 4 + ( 3*3) = 19

Execution time on P2 = 19 / 1.5 = 12.66 ns

P2 is faster.